OpenStudy (experimentx):

Evaluate this integral: \[ \large \int_0^\infty\frac1{x^x}dx \]

OpenStudy (experimentx):



OpenStudy (turingtest):

\[\large \int_0^\infty\frac1{x^x}dx\]learn latex!

OpenStudy (experimentx):

yeah i know ... draw is just too quicker.

OpenStudy (anonymous):

I doubt you can define it. The lower bound is an essential singularity.

OpenStudy (anonymous):

@experimentX can u solve this.

OpenStudy (experimentx):

no not really ... i was expecting it's answer to be in infinite series ... http://openstudy.com/users/leksi#/updates/4ff409aee4b01c7be8c7845f but Mathematica refused to give it's answer. http://www.wolframalpha.com/input/?i=integrate+1%2Fx^x+from+0+to+infinity

OpenStudy (anonymous):

Huh, interesting. Well, I would take a look in Gradshteyn and Ryzhik. If it can be defined at all, even as a series, it will be there.

OpenStudy (experimentx):

i hope so!!

OpenStudy (anonymous):

santosh just like that integral \[ \int_0^\infty\frac1{x^x}dx=\sum_{n=0}^{\infty} \frac{(-1)^n}{n!} \int_{-\infty}^{\infty} t^n e^{(n+1)t} dt\]

OpenStudy (anonymous):

but the later integral part diverges

OpenStudy (experimentx):

yep ... this isn't analytic :((

OpenStudy (anonymous):

try it with mathematica...in terms of n and t

OpenStudy (anonymous):

Hmm, if you plot the integrand it seems perfectly well-behaved over the entire domain.

OpenStudy (experimentx):

@Carl_Pham it converges ... @mukushla that's one hell of crazy structure :(((

OpenStudy (anonymous):

@experimentX please tell me how